OpenStudy (anonymous):

For f(x)= (x-2)^1/5 + 3, calculate the maximum value that delta can be so that f(x) is within 0.2 unit of 3 when x is within delta units of 2.

OpenStudy (anteater):

So your epsilon is 0.2, and your limit is 3\[\left| x - 2 \right| < \delta \rightarrow \left| \sqrt[5]{x-2 }+ 3 - 3\right| < 0.2\]

OpenStudy (anteater):

\[\left| \sqrt[5]{x-2} \right| < 0.2\]

OpenStudy (anteater):

\[\left| x-2 \right| < 0.2^{5}\]

OpenStudy (anteater):

\[\left| x-2 \right| < .00032\]

OpenStudy (anteater):

So, delta = .00032

OpenStudy (anteater):

That is, when x is within .00032 units of 2, f(x) is within 0.2 units of 3

OpenStudy (anteater):

I hope that helps!